# Quasi-homogeneous singularities of projective hypersurfaces and Jacobian syzygies

### v1

@@ 2.1. Jacobian ideal of a hypersurface
It is well known that $d_{1}=0$, that is the $n+1$ partials $\partial_{0}f$, $\partial_{1}f$, $\cdots$, $\partial_{n}f$ are linearly dependent, if and only if $V$ is a cone with vertex of point $p\in\mathbb{P}^{n}$ of multiplicity $d$. In this last case, the point $p$ is the only singular point of $V$ and it is easily seen to be a quasi-homogeneous singular point, since the Jacobian scheme is a complete intersection. Therefore we will always assume that $d_{1}>0$.
@@
    The rational quintic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{0}x_{1}^{2}x_{2}^{2}+x_{1}^{5}+x_{2}^{5})$ is singular only in $p=(1:0:0)$ and $J_{f}=(x_{1}^{2}x_{2}^{2},5x_{1}^{4}+2x_{0}x_{1}x_{2}^{2},2x_{0}x_{1}^{2}x_{2}+5x_{2}^{4})$. We have $\tau(C)=\tau_{p}(C)=10$. By [10, Proposition 4.10] one has $\mu(C)=\mu_{p}(C)=11$ and the singularity is non-quasi-homogeneous.
@@ 3. Main result and several reformulations
see also [9, Section 2]). In other words, we have a relation
@@ 4. A geometric characterization of quasi-homogeneous singularities of reduced plane curves
In this framework, the symmetric algebra $S(J_{f})=\oplus_{n\geq 0}Sym^{n}J_{f}$ of the Jacobian ideal will play an important role. In particular, we will strongly use the Cohen-Macaulayness of $S(J_{f})$ which follows from [18, Corollary 10.4]. In fact, we have:
@@
It follows from [18, Corollary 10.4] taking into account that a graded ring is Cohen-Macaulay if and only if its localization at the maximal homogeneous ideal is Cohen-Macaulay (see [19, Proposition 4.10]). ∎
@@
The following examples of curves with non-quasi-homogeneous singularities are taken from [21, Example 2.6].

### v2

@@ 1. Introduction
Moreover, there is a parallelism between the notions of degree and Hilbert-Samuel multiplicity of a non-reduced zero dimensional connected component of the base scheme of a projective linear system, and the Tjurina and Milnor numbers of an isolated singularity, and being locally a complete intersection translates into a quasi-homogeneous singularity. The study of the difference between the two numbers is considered in [6] by comparing the Rees algebra and the symmetric algebra of the ideal defining a generically finite rational map from $\mathbb{P}^{n}\dasharrow\mathbb{P}^{n+1}$. By [21, Theorem 1], the polar map of a hypersurface $V(f)$ with only isolated singularities is generically finite, unless $V(f)$ is a cone with vertex a point. When the two degrees are different, the authors prove in [6, Proposition 5] that the additional component of the symmetric algebra is a hyperplane of multiplicity equal to the difference of the two degrees. We point out, however, that Theorem 3.1 holds in all cases.
@@ 2.1. Jacobian ideal of a hypersurface
It is well known that $d_{1}=0$, that is the $n+1$ partials $\partial_{0}f$, $\partial_{1}f$, $\cdots$, $\partial_{n}f$ are linearly dependent, if and only if $V$ is a cone with vertex of point $p\in\mathbb{P}^{n}$ of multiplicity $d$. In this last case, the point $p$ is the only singular point of $V$ and it is easily seen to be a quasi-homogeneous singular point, since the Jacobian scheme is a complete intersection. Therefore, we will always assume that $d_{1}>0$.
@@
    The rational quintic $C\subset\mathbb{P}^{2}$ of equation $C=V(f)=V(x_{0}x_{1}^{2}x_{2}^{2}+x_{1}^{5}+x_{2}^{5})$ is singular only in $p=(1:0:0)$ and $J_{f}=(x_{1}^{2}x_{2}^{2},5x_{1}^{4}+2x_{0}x_{1}x_{2}^{2},2x_{0}x_{1}^{2}x_{2}+5x_{2}^{4})$. We have $\tau(C)=\tau_{p}(C)=10$. By [11, Proposition 4.10] one has $\mu(C)=\mu_{p}(C)=11$ and the singularity is non-quasi-homogeneous.
@@ 3. Main result and several reformulations
see also [10, Section 2]). In other words, we have a relation
@@
Tohăneanu has shown in [28, Theorem 2.2] that a curve $C\subset\mathbb{P}^{2}$ having only isolated singularities is free if and only it admits a Jacobian syzygy which is non-zero at any point of $\mathbb{P}^{2}$. The following result tells what happens when the curve is not free, and in fact this property holds in any dimension.
Proposition 3.7.
Let $V:f=0$ be a hypersurface in $\mathbb{P}^{n}$ having only isolated quasi-homogeneous singularities. Then for any countable subset $Y$ in $\mathbb{P}^{n}$, there is a syzygy of $V$ of degree at most $d_{m}$, the maximal exponent of $V$, which is non-zero at any point of the set $Y$. In particular $Y$ can be the singular set of $V$.
Proof.
Consider the finite dimensional $\mathbb{C}$-vector space
$W={\rm Syz}(J_{f})_{d_{m}}.$
For any point $y\in Y$ we denote by $E_{y}$ the vector subspace of $W$ consisting of all the Jacobian syzygies of $V$ in $W$ which vanish at the point $y$. If $y\in Y$ is a singular point of $V$, then Theorem 3.1 in the form given in Corollary 3.4 shows that there is a syzygy in $W$ not vanishing at $y$ and hence we have $E_{y}\neq W$.
If $y\in Y$ is not a singular point of $V$, at least one of the partial derivatives $\partial_{j}f$ does not vanish at $y$. It follows that one of the syzygies $\rho_{ij}$ of degree $(d-1)$, associated to the Koszul relation
$\partial_{j}f\partial_{i}f-\partial_{i}f\partial_{j}f=0$
is non zero at $y$. Note that we have $(d-1)\leq d_{m}$, since $d_{m}$ is the maximal exponent of $V$. Hence in this case we also have $E_{y}\neq W$.
Now applying Baire Theorem or some basic facts of Measure Theory, we see that
$E=\cup_{y\in Y}E_{y}\neq W.$
It follows that any syzygy in $W\setminus E$ satisfies our claim. ∎
@@ 4. A geometric characterization of quasi-homogeneous singularities of reduced plane curves
In this framework, the symmetric algebra $S(J_{f})=\oplus_{n\geq 0}Sym^{n}J_{f}$ of the Jacobian ideal will play an important role. In particular, we will strongly use the Cohen-Macaulayness of $S(J_{f})$ which follows from [19, Corollary 10.4]. In fact, we have:
@@
It follows from [19, Corollary 10.4] taking into account that a graded ring is Cohen-Macaulay if and only if its localization at the maximal homogeneous ideal is Cohen-Macaulay (see [20, Proposition 4.10]). ∎
@@
The following examples of curves with non-quasi-homogeneous singularities are taken from [23, Example 2.6].